OpenStudy (avai70178):

y=2+(x-3) Can anyone put his in slope intercept form?

OpenStudy (anonymous):

put what?

OpenStudy (avai70178):

y=2+(x-3)

OpenStudy (anonymous):

could you please remind me with the slope intercept form?

OpenStudy (anonymous):

Y=mx+b, true?

OpenStudy (avai70178):

yes

OpenStudy (anonymous):

what is your B, or y-int value?

OpenStudy (anonymous):

it is -1, isn't it?

OpenStudy (anonymous):

your m will be equal to 1, which is your slope!

OpenStudy (anonymous):

I think you now have it, Y= (1)X-1 y=x-1

OpenStudy (anonymous):

I think it is y+2x-(-6)

OpenStudy (anonymous):

Sorry I ment y=2x+(-6)

OpenStudy (avai70178):

why did you distribute

OpenStudy (anonymous):

??

OpenStudy (avai70178):

no no not you you helped ty

OpenStudy (avai70178):

the other person did something totally different

OpenStudy (anonymous):

oh ok

OpenStudy (avai70178):

haha yeah

OpenStudy (anonymous):

Because that is how it says to do it in my math book. Why I distributed. you have to get y alone the problem in order.

OpenStudy (anonymous):

So the points would be ( 2, -6) when graphed.